OpenStudy (anonymous):

Factor the four-term polynomial. mn - 15 + 3m - 5n (m + 3)(n - 5) (n - 3)(m - 5) (n + 3)(m - 5)

OpenStudy (anonymous):

I think it is the 1st one

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

okay how do u do it

OpenStudy (whpalmer4):

You should multiply out each of those expressions...you need the practice!

OpenStudy (whpalmer4):

The 1st one: \[(m+3)(n-5) = mn -5m + 3n - 15\]

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

Let (m+x) (n+y) = mn - 15 + 3m - 5n => mn+ym+xn+xy=mn-15+3m-5n

OpenStudy (anonymous):

U need to find x and y from there

OpenStudy (whpalmer4):

\[(a+b)(c+d) = a(c+d) + b(c+d) = ac + ad + bc + bd\] If you want to be clever, note that to get a +3m, you need to have a +3 multiply an m, and there's only one such equation.

OpenStudy (anonymous):

comparing coffecients gives: x=-5 and y=3

OpenStudy (anonymous):

so, (m+x) (n+y) = mn - 15 + 3m - 5n = (m-5)(n+3)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so its the last one

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay lol help me again please